A 1 inch diameter nozzle there is a water flows under 400ft head to operate a turbine. There is 90% efficient of turbine and connected to a 94% efficient generators. Find the power output in kilowatts.


1
Expert's answer
2021-05-19T06:07:50-0400

The water flows through a nozzle of 1 inch = 0.0254 m


head loss 400 ft. = 121.92 m


Area = "\pi r^2" = 3.142 "\times (\frac{0.0254}{2})^2= 5.068 \times 10^{-4} m^2"


Assuming that the flow rate is 3 m/s the mass of water of = 3000 kg/s


Hydropower W = "m \times g \times H_{net} \times \eta"


"\eta =" "0.90 \times 0.94= 0.846 = 84.6" %


assuming the that the head loss = 10%


the net head "H_{net}= 0.9\times" 121.92 = 109.728 m


"\eta = 3000 \times 9.8\times 109.728 \times 0.846= 2.7 Megawatts"